As you are most probably aware, (se previous post on January 12), segment 1 of Arm B had a severe temperature problem starting on December 31st. This was solved early in January but a total recalibration of the system was necessary. This calibration was finsihed a couple of days ago by the L1 team . We are now finishing checking the overall performance figure but we can say already that SMOS is completely back to its previous performance  level (i.e., before  mid December). During the next level 1 reprocessing erxercise, the whole data set will be brought to the same level of calibration.

Also note that later today we will update the link performances and version with the latest results where you will be able to see the current impact of the arm B1 issue.

From today on this page will be updated daily.

Finally ESA has finished reprocessing level1c for the whole year 2010 and at CESBIO we are reprocessing level2 soil moisture. Check this blog to know as soon as it is done and verified!

Leave a Reply

Your email address will not be published. Required fields are marked *

This site uses Akismet to reduce spam. Learn how your comment data is processed.